# Tianxiang Chen USTC, China Alibaba Cloud, China

### v3

@@ 4.1 Experimental Settings
4.1.3 Optimization.
The algorithm is implemented in Pytorch, with Adaptive Gradient (AdaGrad) as the optimizer with the initial learning rate set to 0.05 and weight decay coefficient set to 0.0004. 2 NVIDIA A6000 GPU is used for training, with batch size set to 32. Dice loss [48] is adopted as the loss function. Training on SIRST and IRSTD-1k takes 800 epochs and 600 epochs respectively.
Table 1: Accuracy comparison on NUAA-SIRST and IRSTD-1k. The figures in bold and underline mark the highest and the 2nd highest ones in each column.
@@
IAANet [20] Hybrid 75.31 74.65 98.22 35.65 59.82 58.24 88.62 24.79
RKformer [24] Hybrid 77.24 74.89 99.11 1.580 64.12 64.18 93.27 18.65
@@
TCI-Former [11] Hybrid 80.79 79.85 99.23 4.189 70.14 67.69 96.31 14.81
MiM-ISTD Mamba 80.80 80.20 100 2.179 70.33 67.82 96.91 13.49
4.2 Quantitative Comparison
4.2.1 Accuracy Comparison.
We select some of the SOTA ISTD methods for comparison. As shown in Table 1, our MiM-ISTD generally achieves the best performances in terms of pixel-level metrics and object-level metrics on both datasets.
@@ 4.2 Quantitative Comparison
For the object-level metrics ($P_{d},F_{a}$), how to reach a trade-off between $P_{d}$ and $F_{a}$ is challenging because the two metrics are mutually exclusive: higher $P_{d}$ also increases the possibility of higher $F_{a}$. From Table 1, we can see that our method generally achieves the best object-level metrics results compared with other methods, especially detecting all the small targets ($P_{d}=100\%$) in the NUAA-SIRST test set. The result demonstrates that our MiM-ISTD can learn better representations to overcome missed detection and false alarms.
4.2.2 Efficiency Comparison.
We also compare the efficiency of different methods in terms of parameter number (M), FLOPs (G) and inference time (s) and GPU memory usage (M) during training on $512\times 512$ infrared image datasets, as shown in Table 2. Compared with other methods except ACM, our MiM-ISTD in general has significantly less parameters, GFLOPs, inference time and memory usage. This is because we adopt the efficient Mamba structure and also use a shared network to calculate the relations of each visual words in visual sentences so that the increased parameters and GFLOPs is negligible. Even compared with the most lightweight ISTD model ACM, our MiM-ISTD still has less GFLOPs and significantly higher accuracy. In addition, in higher resolution infrared image scenarios, we have compared in Fig. 1 that MiM-ISTD’s superiority of inference time and GPU memory usage will further be expanded while always maintaining a superior accuracy. Generally, MiM-ISTD reaches the best balance between efficiency and accuracy.
Table 2: Comparison of the model parameters (M), FLOPs (G), inference time (s) per image and GPU memory (M) per 4 batch size of different methods.
@@
MiM-ISTD 1.16 1.01 0.03 1774 74.01
Figure 5: Visualization of predicted mask results.
4.3 Visualization
4.3.1 Visualization of Mask Results.
Visual results with closed-up views of different methods are shown in Fig. 5, where present methods more or less suffer from incomplete detection and missed detection. Compared with other SOTAs, our MiM-ISTD better curtails these cases and more completely detect the shapes of all small targets. This is because integrating an inner Mamba block assists our network to further exploit more local features, which promotes more refined detection of small targets. This can also be proved by comparing MiM-ISTD visual results with the "no inner Mamba" visual results, where abandoning inner Mamba brings worse detection performances.
4.3.2 Visualization of Feature Maps.
We visualize the learned features of TCI-Former and MiM-ISTD to further understand the effect of the proposed method. The feature maps are formed by reshaping the patch embeddings according to their spatial positions. The feature map outputs of stage 1, 2, 3 and 4 are shown in Figure  6 (a) where 9 feature map channels are randomly sampled for each of these outputs. In MiM-ISTD, the local information are better preserved in deeper layers compared to TCI-Former. Also, MiM-ISTD has higher feature consistency among each channels than TCI-Former, meaning that the features extracted by MiM-ISTD are more focused on the target. These benefits owe to the introduction of inner Mamba block to further model local features. We additionally visualize all the 64 channels of the stage 3 feature output using t-SNE [49] in Figure  6 (b) to demonstrate our analysis. We observe that the features of MiM-ISTD are more concentrated, exhibiting stronger discriminative power.
Figure 6: Visualization of the features of TCI-Former and MiM-ISTD.
Table 3: Ablation study of each module on NUAA-SIRST.
@@ 4.3 Visualization
##\##2 Inner Mamba $\rightarrow$ Conv 78.47 76.26 99.15 5.014
##\##3 Ours 80.80 80.20 100 2.179
Table 4: Ablation study of the granularity of patch division on NUAA-SIRST.
@@
##\##1 Visual Word $\rightarrow 1\times 1$ 78.71 76.43 99.12 5.965
##\##2 Visual Word $\rightarrow 4\times 4$ 79.50 77.68 100 2.980
##\##3 Visual Sentence $\rightarrow H/4\times W/4$ 79.67 78.04 100 3.685
##\##4 Visual Sentence $\rightarrow H/16\times W/16$ 69.32 67.58 93.52 22.21
##\##5 Ours (VW $\rightarrow 2\times 2$, VS $\rightarrow H/8\times W/8$ ) 80.80 80.20 100 2.179
4.4 Ablation study
4.4.1 Ablation study of Each Module.
The ablation study of each module is shown in Table 3. Our baseline (No.#$\#1$) uses the visual Mamba [28] encoder as our encoder with descending spatial resolutions, where no inner Mamba block is employed. We also replace the inner Mamba block of our MiM-ISTD with the standard convolution, batch normalization and activation operations (No.#$\#2$) to examine the effect of our inner Mamba block. We find that our present setting brings the best results, because convolution typically focus on local information, while the inner Mamba block can capture complex spatial relationships and integrate multi-directional information to have a more delicate perception and identification of small targets.
4.4.2 Ablation Study of the Granularity of Patch Division.
To explore the impact of information contained in each visual word produced by the convolutional stem, we adjust the representation granularity of sub-patches so that each visual word corresponds to a $1\times 1$ (No.#$\#1$) or $4\times 4$ (No.#$\#2$) pixel region in the original image, in contrast to $2\times 2$ of our present setting. We also change the initial spatial shape of visual sentence from $\frac{H}{8}\times\frac{W}{8}$ to $\frac{H}{4}\times\frac{W}{4}$ (No.#$\#3$) and $\frac{H}{16}\times\frac{W}{16}$ (No.#$\#4$) to ablate the effect of visual sentence granularity. We can observe that too large or too small division granularity cannot bring the most ideal performance. We adopt the configuration that demonstrates the best performance in our present setting.
5 Conclusion

### v4

@@ IV-A Experimental Settings
IV-A3 Optimization.
The algorithm is implemented in Pytorch, with Adaptive Gradient (AdaGrad) as the optimizer with the initial learning rate set to 0.06 and weight decay coefficient set to 0.0004. 2 NVIDIA A6000 GPU is used for training, with batch size set to 32. Dice loss [48] is adopted as the loss function. Training on SIRST and IRSTD-1k takes 1000 epochs and 800 epochs respectively.
TABLE I: Accuracy comparison on NUAA-SIRST and IRSTD-1k. The figures in bold and underlined mark the highest and the 2nd highest ones in each column.
@@
IAANet [20] CNN-ViT 75.31 74.65 98.22 35.65 59.82 58.24 88.62 24.79
RKformer [24] CNN-ViT 77.24 74.89 99.11 1.580 64.12 64.18 93.27 18.65
@@
SegFormer [51] ViT 67.64 66.43 89.92 35.83 60.12 57.23 88.92 38.93
TCI-Former [11] CNN-ViT 80.79 79.85 99.23 4.189 70.14 67.69 96.31 14.81
MiM-ISTD Mamba 80.92 80.13 100 2.168 70.36 68.05 96.95 13.38
IV-B Quantitative Comparison
IV-B1 Accuracy Comparison.
We select some of the SOTA ISTD methods for comparison. As shown in Table I, our MiM-ISTD generally achieves the best performances in terms of pixel-level metrics and object-level metrics on both datasets.
@@ IV-B Quantitative Comparison
For the object-level metrics ($P_{d},F_{a}$), how to reach a trade-off between $P_{d}$ and $F_{a}$ is challenging because higher $P_{d}$ also increases the possibility of higher $F_{a}$. From Table I, we can see that our method generally achieves the best object-level metrics results, especially detecting all the small targets ($P_{d}=100\%$) in the NUAA-SIRST test set. The result demonstrates that our MiM-ISTD can learn better representations to overcome missed detection and false alarms.
IV-B2 Efficiency Comparison.
We also compare the efficiency of different methods in terms of parameter number (M), FLOPs (G) and inference time (s) and GPU memory usage (M) during training on $512\times 512$ infrared image datasets, as shown in Table II. Compared with other methods except ACM, our MiM-ISTD has significantly fewer parameters, GFLOPs, inference time, and memory usage. This is because we adopt the efficient Mamba structure and also use a shared network to calculate the relations of each visual word in visual sentences so that the increased parameters and GFLOPs are negligible. Even compared with the most lightweight ISTD model ACM, our MiM-ISTD still has fewer GFLOPs and significantly higher accuracy. In addition, in higher resolution infrared image scenarios, we show in Fig. 1 that MiM-ISTD’s superiority of inference time and GPU memory usage will further be expanded while still maintaining a superior accuracy. Notably, compared with the version without Inner Mamba blocks, we notice a slight decrease in model complexity, parameter count, and GPU memory usage. However, the inference speed remains unchanged, and there is a notable decline in average accuracy. This suggests that while the model becomes marginally lighter without the inner Mamba blocks, this comes at the expense of its overall accuracy. Generally, MiM-ISTD reaches the best efficiency-accuracy balance.
TABLE II: Comparison of the model parameters (M), FLOPs (G), inference time (s) per image, and GPU memory (M) per 4 batch size of different methods.
@@
w/o Inner Mamba 4.67 3.91 0.03 1434 70.86
MiM-ISTD 4.76 3.95 0.03 1996 74.09
Figure 5: Visualization of predicted mask results. The predicted targets are amplified in the yellow boxes for clearer observation, and we highlight some inaccuracies, such as false alarms and missed detection, made by other methods.
IV-C Visualization
IV-C1 Visualization of Mask Results.
Visual results with closed-up views of different methods are shown in Fig. 5, where present methods more or less suffer from incomplete detection and missed detection. Compared with other SOTAs, our MiM-ISTD better curtails these cases and more completely detects the shapes of all small targets. This is because integrating an Inner Mamba block assists our network to further exploit more local features, which promotes more refined detection of small targets. This can also be proved by comparing MiM-ISTD visual results with the ”no Inner Mamba” visual results, where abandoning the Inner Mamba brings worse detection performances.
IV-C2 Visualization of Feature Maps.
We visualize the learned features of TCI-Former and MiM-ISTD to further understand the effect of the proposed method. The feature maps are formed by reshaping the patch embeddings according to their spatial positions. The feature map outputs of stages 1, 2, 3, and 4 are shown in Fig. 6a, where 9 feature map channels are randomly sampled for each of these outputs. In MiM-ISTD, the local information is better preserved in deeper layers compared to TCI-Former. Also, MiM-ISTD has higher feature consistency among each channel than TCI-Former [11], the present SOTA method, meaning that the features extracted by MiM-ISTD are more focused on the target. These benefits are owed to the introduction of the Inner Mamba block to further model local features. We additionally visualize 64 channels of the stage 3 feature output using t-SNE [52] in Fig. 6b to demonstrate our analysis. We observe that the features of MiM-ISTD are more concentrated than the model without Inner Mamba blocks and the present SOTA method [11]. This observation aligns with the results obtained from feature comparison. In general, our MiM-ISTD exhibits stronger discriminative power.
Figure 6: (a) Feature map outputs of Stage 1/2/3/4. and (b) Visualization through T-SNE of feature outputs from the third stage, across 64 channels, using different methods.
Figure 7: ROC curves on (a) NUAA-SIRST and (b) IRSTD-1k datasets.
IV-D ROC results
While $IoU$, $nIoU$, $P_{d}$ and $F_{a}$ measure the segmentation performance under a fixed threshold, the ROC can provide an overall evaluation under multiple different thresholds. We also compare the ROCs among other SOTA methods in Fig. 7. It can be noted that our MiM-ISTD exhibits overall impressive performance on both datasets, particularly within intervals of low false positive rates where its true positive rate swiftly escalates, demonstrating robust detection capabilities. Additionally, at higher false positive rates, MiM-ISTD maintains a relatively high true positive rate, indicating good overall robustness. The results further demonstrate the superiority of our method over other SOTA methods.
IV-E Ablation study
IV-E1 Ablation study of Each Module.
The ablation study of each module is shown in Table III. Our baseline (No.#$\#1$) uses the plain visual Mamba [28] encoder as our encoder with descending spatial resolutions, where no Inner Mamba block is employed. We also replace the Inner Mamba block of our MiM-ISTD with the standard convolution, batch normalization and activation operations (No.#$\#2$) to examine the effect of our Inner Mamba block. We find that our present setting brings the best results, because convolution typically focus on local information, while the Inner Mamba block can capture complex spatial relationships and integrate multi-directional information to have a more delicate perception and identification of small targets.
TABLE III: Ablation study of each module on NUAA-SIRST.
@@ IV-E Ablation study
[…]
##\##3 Ours 80.92 80.13 100 2.168
TABLE IV: Ablation study of the granularity of patch division on NUAA-SIRST.
@@
##\##1 VW $\rightarrow 1\times 1$ 78.71 76.43 99.12 5.965
##\##2 VW $\rightarrow 4\times 4$ 79.50 77.68 100 2.980
##\##3 VS $\rightarrow 4\times 4$ 79.67 78.04 100 3.685
##\##4 VS $\rightarrow 16\times 16$ 69.32 67.58 93.52 22.21
##\##5 Ours (VW $\rightarrow 2\times 2$, VS $\rightarrow 8\times 8$ ) 80.92 80.13 100 2.168
IV-E2 Ablation Study of the Granularity of Patch Division.
To explore the impact of information contained in each visual word produced from the convolutional stem from the start, we adjust the representation granularity of sub-patches so that each visual word corresponds to a $1\times 1$ (No.#$\#1$) or $4\times 4$ (No.#$\#2$) pixel region in the original image, in contrast to $2\times 2$ of our present setting. We also fix the $2\times 2$ reception field of visual word and change the initial spatial shape of visual sentence from $\frac{H}{8}\times\frac{W}{8}$ to $\frac{H}{4}\times\frac{W}{4}$ (No.#$\#3$) and $\frac{H}{16}\times\frac{W}{16}$ (No.#$\#4$) to ablate the effect of visual sentence granularity. We can observe that too large or too small division granularity cannot bring the most ideal performance. We adopt the configuration that demonstrates the best performance in our present setting.
V Conclusion